OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the second and third equations, then the first and second equations.

OpenStudy (anonymous):

5x + 2y + z = –2 3x + 4y + 3z = 2 –4x – 3y – 3z = 1 A. 5x + 2y = –2 3x + 4y = 2 B. –x + y = 3 –12x – 2y = 8 C. 5x + 2y = –2 –12x – 2y = 8 D. x – y = –3 3x + 4y = 2